OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x + y = 6 x = y + 2 (Points : 4) x + y = y + 2 x + (y + 2) = 6 x = (x 6) + 2 (y + 2) + y = 6 i am confused here

OpenStudy (anonymous):

It should be x+y=6 ........(1) x=y+2 ........(2) put equation 2 in 1 (y+2)+y=6 2y+2=6 so y=2 put in y=2 in equation 2 x=4 there it is

OpenStudy (anonymous):

x + y = y + 2 6=y+2 6-2=y y=4

OpenStudy (anonymous):

thanks gotta go thanks for the help

OpenStudy (anonymous):

x + (y + 2) = 6 x+4+2=6 x+6=6 x=6/6 x=1

OpenStudy (anonymous):

forget my hahaha